# Towards deployment-centric multimodal AI beyond vision and language

### v1

@@ Abstract
Multimodal artificial intelligence (AI) integrates diverse types of data via machine learning to improve understanding, prediction, and decision-making across disciplines such as healthcare, science, and engineering. However, most multimodal AI advances focus on models for vision and language data, while their deployability remains a key challenge. We advocate a deployment-centric workflow that incorporates deployment constraints early to reduce the likelihood of undeployable solutions, complementing data-centric and model-centric approaches. We also emphasise deeper integration across multiple levels of multimodality and multidisciplinary collaboration to significantly broaden the research scope beyond vision and language. To facilitate this approach, we identify common multimodal-AI-specific challenges shared across disciplines and examine three real-world use cases: pandemic response, self-driving car design, and climate change adaptation, drawing expertise from healthcare, social science, engineering, science, sustainability, and finance. By fostering multidisciplinary dialogue and open research practices, our community can accelerate deployment-centric development for broad societal impact.
@@ Introduction
Data drives discovery in the 21st century wang2023scientific. From satellites monitoring climate change to social media capturing human behaviour, the variety and volume of information have never been greater. Each type of data, or modality, offers unique insights, but unimodal approaches often fall short of achieving robust or generalisable performance. Autonomous vehicles relying solely on visual data struggle with object detection in low-light or adverse weather conditions kabir2025terrain. During the COVID-19 pandemic, relying solely on RT-PCR data for diagnosing SARS-CoV-2 infection led to frequent false negatives, hindering timely interventions woloshin2020false. Instead, integrating information from multiple modalities can reveal patterns and solutions that unimodal approaches miss Zhao2024DeepMD.
Multimodal artificial intelligence (AI) leverages multimodal data for better understanding complex systems through machine learning baltruvsaitis2018multimodal; ektefaie2023multimodal; xu2023multimodal; liang2024foundations. Its promise is evident in multidisciplinary applications such as pandemic response. For example, in healthcare, combining medical imaging, genomic sequencing, and epidemiological data can enhance diagnoses, inform treatment strategies, and support disease prevention acosta2022multimodal; kline2022multimodal; krones2025review. In science, fusing genetic and protein structure data holds promise for advancing vaccine development notin2024machine. In engineering, integrating textual specifications with spatial data can improve product design and manufacturing song2024multi, which could extend to optimising ventilator production. Across other disciplines, such as sustainability, finance, and social science, multimodal AI offers the potential to provide deeper insights and actionable strategies ofodile2024predictive; quatrini2021challenges; gupta2021emotion; anshul2023multimodal. However, many current approaches face challenges in deployment due to data limitations, integration complexities, and practical constraints roberts2021common.
While multimodal AI receives increasing attention and holds great promise, research has primarily focused on vision-language models bordes2024introduction; zhang2024vision, leaving other modalities, such as tabular and time-series data, and related disciplines underexplored vanposition (Fig. 1 and Supplementary Fig. S2). Real-world challenges, such as pandemic response, call for AI capable of integrating diverse types of data through multidisciplinary collaboration, bridging the gap between research and application.
@@
Figure 1: Trends in multimodal AI research and dominance of vision and language. a, Yearly growth of multimodal AI preprints on arXiv by modality, showing a steady increase over time and the dominance of vision and language. b, Breakdown of modality pairs in multimodal AI preprints on arXiv in 2024, revealing that over half the studies involve vision and language, followed by vision and others (19.9%), language and others (18.0%), and other modality combinations (11.1%). This analysis highlights the most common pairwise modality combinations and shows that those involving vision or language dominate, reaching 88.9%. For clarity and space efficiency, only modality pairs exceeding 1.5% are annotated. See Supplementary Fig. S2 for a more detailed illustration and analysis of the multimodal AI landscape, including statistics on triple and quadruple modality combinations.
@@
Traditionally, multimodal AI research has been model-centric baltruvsaitis2018multimodal; ektefaie2023multimodal; xu2023multimodal, focusing on developing new models to outperform existing ones on standardised benchmarks and datasets. The rise of generative AI jo2023promise models such as ChatGPT brown2020language; achiam2023gpt resulted in a shift towards data-centric approaches seedat2023navigating; liu2024visual; li2024llava, emphasising data resources and quality for better model performance. However, the gap between the high expectations and limited real-world impacts roberts2021common; paleyes2022challenges indicates a pressing need for deployment-centric approaches that prioritise real-world applicability, user needs, and ethical considerations, ensuring that AI innovations are both novel and practical, with a positive impact. We advocate a deployment-centric workflow for advancing multimodal AI from research to scalable solutions, built on ideas from the machine learning operations lifecycle guidelines kreuzberger2023machine and technology readiness levels for machine learning systems lavin2022technology.
Figure 2: Deployment-centric multimodal AI: workflow and challenges. a, Deployment-centric multimodal AI workflow designed to meet real-world needs through a structured three-stage process covering planning, development, and deployment, with iterative assessments and updates. In particular, the planning stage considers deployment constraints early to ensure alignment with real-world settings and practical needs. This workflow incorporates multidisciplinary collaboration at all stages to ensure discipline-specific knowledge informs and enhances AI system development. Moreover, we consider multiple levels of data multimodality, from data types and subtypes to views and fidelities, as illustrated with a healthcare example. This broader definition of multimodality offers new perspectives and rich options for leveraging the benefits of multimodality. The system development stage has five steps similar to a standard machine learning pipeline, where the data-centric and model-centric approaches seedat2023navigating are indicated in the figure to highlight their differences from the deployment-centric approach. b, Five multimodal-AI-specific challenges shared across multiple disciplines and real-world applications.
Figure 2a illustrates our deployment-centric workflow in three stages: planning, development, and deployment. The planning stage focuses on defining the problem, determining the suitability of multimodal AI over unimodal AI, understanding real-world constraints, and formulating AI tasks grounded in realistic assumptions. The development stage builds the multimodal AI system to learn predictive models from multimodal data, and the deployment stage assesses real-world performance, feeding back insights to improve earlier steps. Guidance from multidisciplinary collaboration (the top right of Fig. 2a) informs the entire workflow, ensuring that technical, ethical, and societal perspectives are integrated into the process for a robust and responsible system.
@@
The planning stage (the three blocks on the top left of Fig. 2a) begins with problem definition, which involves clearly articulating the problem’s scope and objectives, and preliminarily assessing whether incorporating multimodal data can offer meaningful advantages, such as improved predictive performance or deeper insights, over unimodal approaches huang2021makes. Beyond vision (image, video) and language (text), key modalities include audio, numeric time series (e.g. financial data sequences), sensor signals (e.g. wearable physiological measurements), and spatial (geolocation), tabular (structured data such as clinical records), and graph-based (relationships or networks) data. Furthermore, multimodality exists at multiple levels: multiple data types, subtypes, views, and fidelities meng2021multi; Penwarden2022MultifidelityPINNs (the middle right of Fig. 2a). For example, in healthcare, multimodal data can range from different types, e.g. electronic health records (EHRs), medical images, and omics data zitnik2019machine; lunke2023integrated, to subtypes within a data type, e.g. X-ray, magnetic resonance imaging (MRI), and ultrasound within the imaging modality, multiple views of the same data (sub)type, e.g. anteroposterior, lateral, and oblique views of X-ray, and multiple fidelities, e.g. high, low, and adaptive fidelities of anteroposterior X-ray. Under this broad definition, the potential number of modalities can expand significantly as more levels are considered. Understanding these modalities early helps select those suited to the problem’s complexity and capable of offering insights beyond unimodal data.
After establishing the potential benefits of multimodality, we move on to deployment constraints, which involves examining the AI system’s application context, including user needs, data availability, regulatory compliance wu2021medical, ethical considerations, societal impact, and economic trade-offs between high- and low-cost data modalities. This step ensures that the chosen modalities and the resulting multimodal AI system are not only technically feasible but also viable within the intended deployment environments. Understanding these constraints early informs task formulation, helping align the solution with real-world requirements and making it more effective and responsible.
The final step, task formulation, translates the defined problem and constraints into specific AI tasks, specifying the inputs, outputs, and evaluation metrics required to meet the objectives set during problem definition. This provides a clear development roadmap, ensuring the multimodal AI system is well-positioned to meet both technical and practical needs.
@@
Developing multimodal AI systems (the lower left of Fig. 2a) parallels standard machine learning system development but introduces added complexities due to the integration of diverse data modalities. We have identified five multimodal-AI-specific challenges (Fig. 2b). Modality incompleteness occurs when one or more modalities are missing during training or deployment, necessitating the development of highly flexible or generative models. Multimodal heterogeneity deals with varying formats, scales, and data structures, necessitating careful integration strategies to ensure interoperability. Cross-modality alignment ensures that the timing and/or meaning of data from different sources are correctly aligned for coherency and consistency, whether temporal (e.g. matching timestamps) or semantic (e.g. identifying data representing the same concept across modalities). Modality complementarity ensures that different modalities contribute complementary information to enhance performance, as more modalities may not improve results, e.g. if they add noise or redundancy. Finally, multimodal privacy risk zhao2024survey; pranjal2023toward arises when independently anonymised datasets become identifiable again (re-identification) through their fusion, thereby revealing sensitive information and necessitating robust privacy-preserving techniques.
@@
Data collection and curation build on the modality choices identified during planning, gathering relevant data from multiple sources and preparing it for AI model training. Diverse and reliable data sources ensure a high-quality, comprehensive representation of the problem space. Once collected, the data undergoes standard curation, such as wrangling, cleaning, annotation, handling missing values, and quality assurance paleyes2022challenges, tailored to the needs of multimodal AI. For instance, in healthcare, linking multimodal data from EHRs, imaging, and biosignals poses significant challenges due to privacy concerns zhao2024survey and the need for standardisation. The integration and cross-referencing of these modalities can amplify the risk of re-identification, underscoring the importance of addressing multimodal privacy risk through robust privacy-preserving techniques al2019privacy; Wendland:2022vo; che2023MultimodalFL. Moreover, selected modalities should be interoperable and complementary, with their alignment and integration resulting in a cohesive, high-quality dataset suited to both technical and deployment requirements.
@@
Multimodal learning integrates diverse data modalities to leverage their unique strengths for improved predictive performance baltruvsaitis2018multimodal. Traditional fusion strategies, early fusion (combining raw features), intermediate fusion (merging processed features), and late fusion (integrating outputs from independently processed modalities), offer trade-offs between model complexity and the timing of modality interaction. While these methods are well-established, recent advances such as hybrid fusion Zhao2024DeepMD and knowledge distillation Wang2020MultimodalLW provide greater flexibility, enabling the combination of multiple strategies or the transfer of knowledge from complex to simpler models. Techniques such as co-attention mechanisms Yu2017MultimodalFB can further enhance integration and performance by dynamically adapting to the context of cross-modal interactions. The choice of strategy depends on the specific problem and data characteristics, balancing integration benefits with the added complexity of managing multiple modalities.
@@
Evaluation and interpretation support the reliability, effectiveness, and fairness of multimodal AI systems, especially in complex, high-stakes environments. Beyond standard performance metrics such as accuracy, it is important to assess the contribution of individual modalities, the alignment and synergy between them, and the system’s robustness to noise and variability. For example, evaluating safety in multimodal models weidinger2024holistic is key for understanding how biases across modalities may interact or amplify one another luccioni2024stable, potentially undermining system reliability. Interpretation murdoch2019definitions; rudin2019stop ensures that model decisions are understandable and transparent. Techniques such as heat maps, $t$-SNE, and decision trees help to visualise and explain how different modalities interact and contribute to outcomes, fostering user trust and supporting error analysis and model refinement.
@@
Deploying multimodal AI systems (the bottom of Fig. 2a) requires infrastructure for hosting, scaling, and ongoing management, with continuous monitoring to ensure the system remains effective and relevant paleyes2022challenges. Unlike unimodal AI, these systems face unique challenges, such as integrating diverse data modalities in real-time environments and maintaining performance in the face of modality-specific issues, such as sensor failures or data stream interruptions. Robust monitoring mechanisms enable the system to detect and compensate for failures in one modality by leveraging information from others. Additionally, real-time multimodal data fusion, where multimodal data may arrive asynchronously, requires careful infrastructure design and adaptive algorithms.
@@
Figure 3: Deployment-centric multimodal AI workflow for the pandemic response use case. This example illustrates how the proposed deployment-centric multimodal AI workflow can better address real-world challenges in pandemic response, drawing on multidisciplinary collaboration, as indicated by the different colours across stages. The process begins with defining the problem of interest properly and carefully considering the space of various deployment constraints in order to formulate specific tasks. Two tasks are shown in the figure: one with a focus on healthcare and the other on social science. Next, we develop multimodal AI systems using diverse data sources supported by multidisciplinary collaboration. Finally, we proceed to real-world deployment and assessment, exemplified by the application of patient surge management.
@@
Pandemic response (Fig. 3) presents a complex challenge requiring coordinated efforts across healthcare and other disciplines khojasteh2022climate. Defining the problem involves assessing early how multimodal AI can offer advantages over unimodal approaches by integrating diverse data sources for deeper insights and more accurate predictions. For example, in social science, combining textual content from surveys or social networks with tabular demographic data can help understand mental health impacts and detect changes in online behaviour elmer2020students; valdez2020social. In sustainability, analysing time-series environmental data, satellite imagery, and sustainability reports can reveal the environmental effects of pandemics liu2020near; zheng2020satellite. In finance, integrating economic indicators, transaction patterns, and market response data with health data can predict socio-economic impacts and inform public health strategies mosser2020central; nicola2020socio. A well-defined problem ensures clear constraints and tasks can be established.
@@
The deployment of multimodal AI in pandemic response faces privacy, technical, and economic constraints. Privacy concerns arise from integrating diverse data sources such as EHRs, social media, and genomic information, which increases the risk of re-identification and amplifies the potential impact of data breaches. Maintaining public trust requires robust privacy-preserving techniques al2019privacy; Wendland:2022vo; che2023MultimodalFL and compliance with regulatory frameworks. Technical challenges include real-time monitoring, where multimodal AI systems must integrate and process data streams rapidly for timely interventions. The heterogeneity of health and social data complicates alignment and standardisation, particularly in international and/or multi-centre collaborations. Achieving minimum predictive accuracy thresholds for outbreak detection and public health decision-making helps build trust in AI-driven systems. Economic constraints, including computational costs and budget limitations, restrict accessibility and scalability. High-cost modalities (e.g. MRI) may be impractical in low-resource settings, so multimodal AI systems should flexibly support lower-cost alternatives (e.g. wearables) to promote equitable deployment.
@@
Task formulation translates the defined problem into specific AI objectives, detailing inputs, outputs, and evaluation metrics for pandemic-related challenges. A primary task is predicting disease outbreaks by integrating multimodal data such as epidemiological records, EHRs, and environmental factors to inform early interventions and resource allocation. Another task is monitoring public adherence to health guidelines using social media data, geospatial information, and biosignals, providing actionable insights to refine policies. Additionally, multimodal AI can support real-time decision-making in healthcare by combining patient data streams, including wearable biosignals, medical images, and clinical notes, for managing patient surges effectively ding2020wearable; charlton_wearable_2022. Task formulation ensures that multimodal AI systems align with technical and practical needs, strengthening pandemic response through actionable insights and tailored interventions.
@@
The development of multimodal AI systems for pandemic response adapts a standard workflow of the five key steps outlined above. Data collection assembles diverse sources, such as EHRs, genomic data, and biosignals, to create comprehensive datasets for outbreak prediction and patient management. During curation, standardisation and alignment address challenges such as integrating time-sensitive data streams from biosignals and geospatial sources. Multimodal learning can employ self-supervised zong2024self and cross-domain yang2014cross techniques to capture cross-modal relationships effectively and integrate domain-specific knowledge. Evaluation requires multi-metric, multi-centre validation han2024randomised; plana2022randomized, ensuring robustness for tasks such as outbreak prediction and public adherence monitoring. Interpretability and explainability enable healthcare professionals and policymakers to trust and act on multimodal AI insights in high-stakes scenarios imrie2023multiple. Tailoring these development steps to pandemic-specific needs ensures adaptability, reliability, and ethical compliance in rapidly changing contexts.

### v2

@@ Abstract
Multimodal artificial intelligence (AI) integrates diverse types of data via machine learning to improve understanding, prediction, and decision-making across disciplines such as healthcare, science, and engineering. However, most multimodal AI advances focus on models for vision and language data, while their deployability remains a key challenge. We advocate a deployment-centric workflow that incorporates deployment constraints early to reduce the likelihood of undeployable solutions, complementing data-centric and model-centric approaches. We also emphasise deeper integration across multiple levels of multimodality through stakeholder engagement and interdisciplinary collaboration to broaden the research scope beyond vision and language. To facilitate this approach, we identify common multimodal-AI-specific challenges shared across disciplines and examine three real-world use cases: pandemic response, self-driving car design, and climate change adaptation, drawing expertise from healthcare, social science, engineering, science, sustainability, and finance. By fostering interdisciplinary dialogue and open research practices, our community can accelerate deployment-centric development for broad societal impact.
Figure 1: Trends in multimodal AI research (2019–2024) and the dominance of vision and language. a, Yearly growth of multimodal AI preprints on arXiv by modality, showing a steady increase over time and the dominance of vision and language. b, Breakdown of modality pairs in multimodal AI preprints on arXiv in 2024, revealing that over half the studies involve vision and language, followed by vision and others (19.9%), language and others (18.0%), and other modality combinations (11.1%). This analysis highlights the most common pairwise modality combinations and shows that those involving vision or language dominate, reaching 88.9%. For clarity and space efficiency, only modality pairs exceeding 1.5% are annotated. See Supplementary Section S2 for methodology and details of the trend analysis. See Supplementary Fig. 1 for a more detailed illustration and analysis of the multimodal AI landscape, including statistics on triple and quadruple modality combinations.
@@ Introduction
Data drives discovery in the 21st century Wang et al. 2023. From satellites monitoring climate change to social media capturing human behaviour, the variety and volume of information have never been greater. Each type of data, or modality, offers unique insights, but unimodal approaches often fall short of achieving robust or generalisable performance. Autonomous vehicles relying solely on visual data struggle with object detection in low-light or adverse weather conditions Kabir et al. 2025. During the COVID-19 pandemic, relying solely on RT-PCR data for diagnosing SARS-CoV-2 infection led to frequent false negatives, hindering timely interventions Woloshin et al. 2020. Instead, integrating information from multiple modalities can reveal patterns and solutions that unimodal approaches miss Zhao et al. 2024a.
Multimodal artificial intelligence (AI) leverages multimodal data for better understanding complex systems through machine learning Baltrušaitis et al. 2019; Ektefaie et al. 2023; Xu et al. 2023; Liang et al. 2024b. Its promise is evident in multidisciplinary applications such as pandemic response. For example, in healthcare, combining medical imaging, genomic sequencing, and epidemiological data can enhance diagnoses, inform treatment strategies, and support disease prevention Acosta et al. 2022; Kline et al. 2022; Krones et al. 2025. In science, fusing genetic and protein structure data holds promise for advancing vaccine development Notin et al. 2024. In engineering, integrating textual specifications with spatial data can improve product design and manufacturing Song et al. 2024, which could extend to optimising ventilator production. Across other disciplines, such as sustainability, finance, and social science, multimodal AI offers the potential to provide deeper insights and actionable strategies Ofodile et al. 2024; Quatrini 2021; Gupta et al. 2021; Anshul et al. 2023.
While multimodal AI receives increasing attention and holds great promise, research has primarily focused on vision-language models Bordes et al. 2024; Zhang et al. 2024 (Fig. 1 and Supplementary Fig. 1), leaving other modalities–such as tabular and time-series data–and related disciplines underexplored van Breugel and van der Schaar 2024 (Fig. 2). Real-world challenges, such as pandemic response, call for AI capable of integrating diverse types of data through interdisciplinary collaboration, bridging the gap between research and application.
As modality integration broadens, multimodal-AI-specific deployment challenges increase, including missing modality Ma et al. 2021; Wu et al. 2024, cross-modal alignment Wang et al. 2022; Liang et al. 2024b, and multimodal privacy risk Zhao et al. 2024b; Pranjal et al. 2023. These issues are compounded by broader barriers such as data limitations, integration complexity, and domain-specific constraints Paleyes et al. 2022. For example, rural healthcare may face limited compute infrastructure; financial services may face regulatory delays; and real-time applications such as autonomous driving demand strict latency control. Multimodal setups often amplify these challenges due to increased system complexity. Simply combining diverse modalities is not enough; real-world success requires proactive alignment with deployment constraints from the outset.
@@
Figure 2: Underexplored modality combinations in multimodal AI (2019–2024). Heatmap of combinations of non-vision, non-language modalities in multimodal AI preprints by year. Each row represents a modality combination, and each column corresponds to a publication year. Darker shades indicate higher counts, with rows ordered by their 2024 totals. The coloured circles on the left identify the modality combinations. For example, in 2023, only two preprints used graph and tabular data (sixth row, fifth column). Time series and sensor data form the most common combination, likely because sensor data are often recorded as time series. Time series and spatial data are the second most common, possibly due to the importance of spatiotemporal modelling. In contrast, combinations involving graph, audio, and tabular data remain sparsely studied. These gaps highlight untapped potential for multimodal AI beyond vision and language. See Supplementary Section S2 for details on data processing and modality extraction.
Figure 3: Deployment-centric multimodal AI: workflow and challenges. a, Deployment-centric multimodal AI workflow designed to meet real-world needs through a structured three-stage process covering planning, development, and deployment, with iterative assessments and updates. In particular, the planning stage considers deployment constraints early to ensure alignment with real-world settings and practical needs. This workflow incorporates stakeholder engagement and interdisciplinary collaboration at all stages to ensure that real-world needs and discipline-specific knowledge inform and enhance AI system development. Moreover, we consider multiple levels of data multimodality, from data types and subtypes to views and fidelities, as illustrated with a healthcare example. This broader definition of multimodality offers new perspectives and rich options for leveraging the benefits of multimodality. The system development stage has five steps similar to a standard machine learning pipeline, where the data-centric and model-centric approaches Seedat et al. 2023 are indicated in the figure to highlight their differences from the deployment-centric approach. b, Five multimodal-AI-specific challenges shared across multiple disciplines and real-world applications: modality incompleteness, where one or more modalities are missing at training or deployment; multimodal heterogeneity, reflecting incompatible formats and data structures; cross-modality alignment, which requires synchronising data in time or meaning; modality complementarity, where the goal is to maximise synergy without introducing redundancy; and multimodal privacy risk, where data fusion increases the chance of sensitive re-identification.
@@
Traditionally, multimodal AI research has been model-centric Baltrušaitis et al. 2019; Ektefaie et al. 2023; Xu et al. 2023, focusing on developing new models to outperform existing ones on standardised benchmarks and datasets. The rise of generative AI Jo 2023 models such as ChatGPT Brown et al. 2020; Achiam et al. 2023 resulted in a shift towards data-centric approaches Seedat et al. 2023; Liu et al. 2023; Li et al. 2023a, emphasising data resources and quality for better performance. However, the gap between high expectations and limited real-world impacts Roberts et al. 2021; Paleyes et al. 2022 indicates a pressing need for deployment-centric approaches that prioritise real-world applicability, user needs, and ethical considerations, ensuring that AI innovations are both novel and practical, with a positive impact. We advocate a deployment-centric workflow for advancing multimodal AI from research to scalable solutions, built on ideas from machine learning operations lifecycle guidelines Kreuzberger et al. 2023 and technology readiness levels for machine learning systems Lavin et al. 2022.
Figure 3a illustrates our deployment-centric workflow in three stages: planning, development, and deployment. Planning focuses on defining the problem, determining the suitability of multimodal AI over unimodal AI, understanding real-world constraints, and formulating AI tasks grounded in realistic assumptions. Development builds the multimodal AI system to learn predictive models from multimodal data, and deployment assesses real-world performance, feeding back insights to improve earlier steps. Guidance from stakeholder engagement and interdisciplinary collaboration (the top right of Fig. 3a) informs the entire workflow, ensuring that the perspectives of domain experts, end-users, and decision-makers are integrated for a robust and socially aligned system Nielsen et al. 2025; Lekadir et al. 2025.
@@
Planning (the three blocks on the top left of Fig. 3a) begins with problem definition, which involves clearly articulating the problem’s scope and objectives, and preliminarily assessing whether incorporating multimodal data can offer meaningful advantages, such as improved predictive performance or deeper insights, over unimodal approaches Huang et al. 2021. Beyond vision (image, video) and language (text), key modalities include audio, numeric time series (e.g. financial data sequences), sensor signals (e.g. wearable physiological measurements), and spatial (geolocation), tabular (structured data such as clinical records), and graph-based (relationships or networks) data. Furthermore, multimodality exists at multiple levels: multiple data types, subtypes, views, and fidelities Meng et al. 2021; Penwarden et al. 2022 (the middle right of Fig. 3a and Fig. 5). For example, in healthcare, multimodal data can range from different types, e.g. medical images, electronic health records (EHRs), and omics data Zitnik et al. 2019; Lunke et al. 2023, to subtypes within a data type, e.g. X-ray, magnetic resonance imaging (MRI), and ultrasound within the imaging modality, multiple views of the same data (sub)type, e.g. anteroposterior, lateral, and oblique views of X-ray, and multiple fidelities, e.g. high, low, and adaptive fidelities of anteroposterior X-ray. Under this broad definition, the potential number of modalities can expand substantially as more levels are considered. Understanding these modalities early helps select those suited to the problem’s complexity and capable of offering insights beyond unimodal data.
After establishing the potential benefits of multimodality, we move on to deployment constraints, which involves examining the AI system’s application context, including user needs, data availability, regulatory compliance Wu et al. 2021, ethical considerations, societal impact, and economic trade-offs between high-cost and low-cost data modalities. This step ensures that the chosen modalities and the resulting multimodal AI system are not only technically feasible but also viable within the intended deployment environments. Understanding these constraints early informs task formulation, helping align solutions with real-world requirements and making it more effective and responsible.
The final step, task formulation, translates the defined problem and constraints into specific AI tasks, specifying the inputs, outputs, and evaluation metrics required to meet the objectives set during problem definition. This provides a clear development roadmap, ensuring multimodal AI systems are well-positioned to meet both technical and practical needs. Selecting modalities requires balancing utility, acquisition cost, complexity, and deployment feasibility. Fewer well-curated modalities may outperform broader but less practical combinations.
@@
Developing multimodal AI systems (the lower left of Fig. 3a) parallels standard machine learning system development but introduces added complexities due to the integration of diverse data modalities. We have identified five multimodal-AI-specific challenges (Fig. 3b). Modality incompleteness occurs when one or more modalities are missing during training or deployment, necessitating the development of highly flexible or generative models. Multimodal heterogeneity deals with varying formats, scales, and data structures, necessitating careful integration strategies to ensure interoperability. Cross-modality alignment ensures that the timing or meaning of data from different sources is correctly aligned for coherency and consistency, whether temporal (e.g. matching timestamps) or semantic (e.g. identifying data representing the same concept across modalities). Modality complementarity ensures that different modalities contribute complementary information to enhance performance, as more modalities may not improve results, e.g. if they add noise or redundancy. Finally, multimodal privacy risk Zhao et al. 2024b; Pranjal et al. 2023 arises when independently anonymised datasets become identifiable again (re-identification) through their fusion, thereby revealing sensitive information and necessitating robust privacy-preserving techniques.
@@
Data collection and curation build on the modality choices identified during planning, gathering relevant data from multiple sources and preparing them for AI model training. Diverse and reliable data sources ensure a high-quality, comprehensive representation of the problem space. Synthetic data and weak supervision provide valuable alternatives when labelled data is scarce van Breugel et al. 2024, such as in rare diseases or time-sensitive crises, where expert annotation may be infeasible. Once collected, the data undergoes standard curation, such as wrangling, cleaning, annotation, handling missing values, and quality assurance Paleyes et al. 2022, tailored to multimodal AI. For instance, in healthcare, linking multimodal data from EHRs, imaging, and biosignals poses substantial challenges due to privacy concerns Zhao et al. 2024b and the need for standardisation. The integration and cross-referencing of these modalities can amplify the risk of re-identification, underscoring the importance of addressing multimodal privacy risk through robust privacy-preserving techniques Al-Rubaie and Chang 2019; Wendland et al. 2022; Che et al. 2023. Moreover, selected modalities should be interoperable and complementary, with their alignment and integration resulting in a cohesive, high-quality dataset suited to technical and deployment requirements.
@@
Multimodal learning integrates diverse data modalities to leverage their unique strengths for improved predictive performance Baltrušaitis et al. 2019. Traditional fusion strategies, early fusion (combining raw features), intermediate fusion (merging processed features), and late fusion (integrating outputs from independently processed modalities), offer trade-offs between model complexity and when modalities interact. Recent advances, hybrid fusion Zhao et al. 2024a and knowledge distillation Wang et al. 2020, provide greater flexibility, enabling the combination of multiple strategies or the transfer of knowledge from complex to simpler models. Techniques such as co-attention mechanisms Yu et al. 2017 can further enhance integration and performance by dynamically adapting to cross-modal interactions. The choice of strategy depends on the specific problem and data characteristics, balancing integration benefits with the added complexity of managing multiple modalities.
@@
Evaluation and interpretation support the reliability, effectiveness, and fairness of multimodal AI systems, especially in complex, high-stakes environments. Beyond standard performance metrics such as accuracy, evaluation should assess the contribution of individual modalities, the alignment and synergy between them, and the system’s robustness to noise and variability. For example, evaluating safety in multimodal models Weidinger et al. 2024 is key for understanding how biases across modalities may interact or amplify one another Luccioni et al. 2023, potentially undermining system reliability. Interpretation Murdoch et al. 2019; Rudin 2019 ensures that model decisions are understandable and transparent. Techniques such as heat maps, $t$-SNE, and decision trees help visualise and explain how different modalities interact and contribute to outcomes, fostering user trust and supporting error analysis and model refinement. Uncertainty estimation Gawlikowski et al. 2023, via ensembling, calibration, or Bayesian approaches, supports more reliable decisions in high-stakes applications.
@@
Deploying multimodal AI systems (the bottom of Fig. 3a) requires infrastructure for hosting, scaling, and ongoing management, with continuous monitoring to ensure the system remains effective and relevant Paleyes et al. 2022. Infrastructural and regulatory conditions, including compute availability, connectivity, and legal safeguards, can determine what is practically feasible and where deployment may succeed or fail. Unlike unimodal AI, these systems face unique challenges, such as integrating diverse data modalities in real-time environments and maintaining performance despite modality-specific issues, such as sensor failures or data stream interruptions. Robust monitoring mechanisms enable the system to detect and compensate for failures in one modality by leveraging information from others. Additionally, real-time multimodal data fusion, where multimodal data may arrive asynchronously, requires careful infrastructure design and adaptive algorithms.
@@
Figure 4: Deployment-centric multimodal AI workflow for three use cases. These examples demonstrate the versatility of the proposed workflow, showing how deployment-centric considerations can be applied across various disciplines to solve complex challenges. a, This example illustrates how the proposed deployment-centric multimodal AI workflow can better address real-world challenges in pandemic response, drawing on stakeholder engagement and interdisciplinary collaboration, as indicated by the different colours across stages. The process begins with defining the problem of interest properly and carefully considering the space of various deployment constraints in order to formulate specific tasks. Two tasks are shown in the figure: one with a focus on healthcare and the other on social science. Next, we develop multimodal AI systems using diverse data sources. Finally, we proceed to real-world deployment and assessment, exemplified by the application of patient surge management. b, In self-driving car design, the workflow incorporates deployment constraints from social science, engineering, sustainability, and finance disciplines. Two key tasks focus on engineering and science considerations, respectively. Multimodal AI development uses multiple data sources, including remote sensing, weather, and materials data, aiming for real-world deployment exemplified by automated intersection turning. c, In climate change adaptation, the workflow considers constraints such as mass deployment cost barriers, minimum data sampling density, and prioritisation of vulnerable regions. Task formulation and multimodal AI development focus on environmental sustainability and financial considerations, reflecting how task and model development are shaped by discipline-specific needs. Early flood warning serves as an example of real-world deployment and assessment.
@@
Pandemic response (Fig. 4a) presents a complex challenge requiring coordinated efforts across healthcare and other disciplines Khojasteh et al. 2022. Defining the problem involves assessing early how multimodal AI can offer advantages over unimodal approaches by integrating diverse data sources for deeper insights and more accurate predictions. For example, in social science, combining textual content from surveys or social networks with tabular demographic data can help understand mental health impacts and detect changes in online behaviour Elmer et al. 2020; Valdez et al. 2020. In sustainability, analysing time-series environmental data, satellite imagery, and sustainability reports can reveal the environmental effects of pandemics Liu et al. 2020; Zheng et al. 2020. In finance, integrating economic indicators, transaction patterns, and market response data with health data can predict socio-economic impacts and inform public health strategies Mosser 2020; Nicola et al. 2020. A well-defined problem ensures clear constraints and tasks can be established.
@@
The deployment of multimodal AI in pandemic response faces privacy, technical, and economic constraints. Privacy concerns arise from integrating diverse data sources such as EHRs, social media, and genomic information, which increases the risk of re-identification and amplifies the potential impact of data breaches. Maintaining public trust requires robust privacy-preserving techniques Al-Rubaie and Chang 2019; Wendland et al. 2022; Che et al. 2023 and compliance with regulatory frameworks. Technical challenges include real-time monitoring, where multimodal AI systems must integrate and process data streams rapidly for timely interventions. The heterogeneity of health and social data complicates alignment and standardisation, particularly in international or multi-centre collaborations. Achieving minimum predictive accuracy thresholds for outbreak detection and public health decision-making helps build trust in AI-driven systems. Economic constraints, including computational costs and budget limitations, restrict accessibility and scalability. High-cost modalities (e.g. MRI) may be impractical in low-resource settings, so multimodal AI systems should flexibly support lower-cost alternatives (e.g. wearables) to promote equitable deployment.
@@
Task formulation translates the defined problem into specific AI objectives, detailing inputs, outputs, and evaluation metrics for pandemic-related challenges. A primary task is predicting disease outbreaks by integrating multimodal data such as epidemiological records, EHRs, and environmental factors to inform early interventions and resource allocation. Another task is monitoring public adherence to health guidelines using social media data, geospatial information, and biosignals, providing actionable insights to refine policies. Additionally, multimodal AI can support real-time decision-making in healthcare by combining patient data streams, including wearable biosignals, medical images, and clinical notes, for managing patient surges effectively Ding et al. 2020; Charlton et al. 2022. Task formulation ensures that multimodal AI systems align with technical and practical needs, strengthening pandemic response through actionable insights and tailored interventions.
@@
Developing multimodal AI systems for pandemic response follows the standard workflow of the five key steps outlined above. Data collection assembles diverse sources, such as EHRs, genomic data, and biosignals, to create comprehensive datasets for outbreak prediction and patient management. During curation, standardisation and alignment address challenges such as integrating time-sensitive data streams from biosignals and geospatial sources. Multimodal learning can employ self-supervised Zong et al. 2024 and cross-domain Yang et al. 2014 techniques to capture cross-modal relationships effectively and integrate domain-specific knowledge. Evaluation requires multi-metric, multi-centre validation Han et al. 2024; Plana et al. 2022, ensuring robustness for tasks such as outbreak prediction and public adherence monitoring. Interpretability and explainability enable healthcare professionals and policymakers to trust and act on multimodal AI insights in high-stakes scenarios Imrie et al. 2023. Tailoring these development steps to pandemic-specific needs ensures adaptability, reliability, and ethical compliance in rapidly changing contexts.